# Towards Generalisable Time Series Understanding Across Domains

### v1

@@ 3.2 Pre-Training on Multi-Domain Time Series
We pre-train our model using masked data modelling (MDM) (He et al., 2022) to learn generalisable time series features across domains. We mask a subset of input tokens and only encode the visible (i.e. non-masked) tokens using an encoder $f(\cdot)$. Afterwards, we complement the encoded tokens with learnable mask tokens and feed them to a decoder $g(\cdot)$, reconstructing the original input tokens.
More precisely, we draw a binary mask ${\bm{m}}\in\{0,1\}^{V_{S}\cdot T^{\prime}}$, following the dual masking strategy proposed in Section 3.4, and apply it to the input sequence ${\bm{E}}\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$. Thus, we obtain a visible view ${\bm{E}}[\displaystyle{\bm{m}}]\in\mathbb{R}^{N_{1}\times D}$, where $N_{1}=\sum_{v=1}^{V_{S}}\sum_{t=1}^{T^{\prime}}m_{v,t}$ and $N_{0}=(V_{S}\cdot T^{\prime})-N_{1}$ denote the number of visible and masked tokens, respectively. The visible view ${\bm{E}}[\displaystyle{\bm{m}}]$ is then fed to the encoder $f(\cdot)$ to compute the token features ${\bm{H}}\in\mathbb{R}^{N_{1}\times D}$:
${\bm{H}}=f(\displaystyle{\bm{E}}[\displaystyle{\bm{m}}]).$ (1)
To reconstruct the original input, these token features are fed to the decoder $g(\cdot)$ together with a special, learnable mask token ${\bm{e}}^{\mathcal{M}}\in\mathbb{R}^{1\times D}$, that is inserted at the masked positions where $m_{v,t}=0$:
${\bm{h}}^{\prime}_{v,t}=\begin{cases}\displaystyle{\bm{h}}_{v,t}&\text{if }m_{v,t}=1\\ \displaystyle{\bm{e}}^{\mathcal{M}}&\text{if }m_{v,t}=0\end{cases},$ (2)
such that ${\bm{H}}^{\prime}\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$. The decoder $g(\cdot)$ then predicts the reconstructed input $\widehat{{\bm{X}}}\in\mathbb{R}^{V_{S}\times(T^{\prime}\cdot P)}$:
$\widehat{{\bm{X}}}=g(\displaystyle{\bm{H}}^{\prime}),$ (3)
@@
$\mathcal{L}_{\text{MSE}}=\frac{1}{V_{S}\cdot T^{\prime}}\sum_{v=1}^{V_{S}}\sum_{t=1}^{T^{\prime}}\|\displaystyle{\bm{x}}_{v,t}-\displaystyle\widehat{{\bm{x}}}_{v,t}\|_{2}^{2}.$ (4)
@@ 3.3 Normalised Cross-Correlation Loss
MDM focuses on reconstructing masked parts of the data, emphasising local patterns through the MSE loss (4). However, time series often exhibit long-range dependencies, where past values influence future outcomes over extended periods. To accurately capture these global patterns, we introduce normalised cross-correlation (NCC) as a loss term in MDM for time series:
$\mathcal{L}_{\text{NCC}}=\frac{1}{V_{S}\cdot\overline{T}}\sum^{V_{S}}_{v=1}\sum^{\overline{T}}_{t=1}\hskip 2.84526pt\frac{1}{\sigma_{\displaystyle{\bm{x}}_{v}}\sigma_{\displaystyle\widehat{{\bm{x}}}_{v}}}(x_{v,t}-\mu_{\displaystyle{\bm{x}}_{v}})(\widehat{x}_{v,t}-\mu_{\displaystyle\widehat{{\bm{x}}}_{v}})\in\left[-1,1\right],$ (5)
where $\mu$ and $\sigma$ denote the mean and standard deviation, respectively. Hence, to capture both local and global temporal dynamics, the total loss used to optimise OTiS is defined as
@@
where $\lambda$ is empirically set to $0.1$ during pre-training.
@@ 3.4 Dual Masking Strategy
We design the masking strategy to enhance foundational model capabilites in time series analysis. Specifically, we randomly select between two masking schemes during pre-training, namely random masking and post-fix masking. In 757575\,75of cases, we apply random masking, where each $m_{v,t}$ is independently sampled from a Bernoulli distribution with probability $p=1-\rho$, with $\rho$ denoting the masking ratio (i.e. $m_{v,t}\sim\text{Bernoulli}(1-\rho)$). This encourages the model to learn complex inter-variate relationships across the entire time series. In the remaining $25\,\%$ of cases, we employ post-fix masking, which masks the second half of the temporal dimension, leaving only the first half visible (i.e. $m_{v,t}=\mathbbm{1}_{[t\leq T^{\prime}/2]}$). The prediction of future values solely based on past observations simulates real-world forecasting conditions, helping the model to capture temporal causality. Overall, this dual masking strategy enables OTiS to learn both bidirectional relationships and temporal causality, which are essential for general time series analysis.
3.5 Fine-Tuning & Inference on (Unseen) Target Domains
Inclusion of Unseen Domains.
For a new domain $S$, a randomly initialised variate embedding ${\bm{E}}^{\mathcal{V}}_{\mathcal{S}}\in\mathbb{R}^{V_{S}\times D}$ is introduced. The domain-specific tokeniser is then fine-tuned alongside the encoder $f(\cdot)$, and, if required, the decoder $g(\cdot)$, for the specific downstream task, as described in the following.
@@ 3.5 Fine-Tuning & Inference on (Unseen) Target Domains
We use the encoder $f(\cdot)$ and the unmasked input sequence ${\bm{E}}$ to compute all token features ${\bm{H}}=f(\displaystyle{\bm{E}})\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$. We average-pool these features into a global token ${\bm{h}}^{*}\in\mathbb{R}^{1\times D}$, which we feed through a linear layer to obtain the final model prediction. We optimise a cross-entropy and MSE loss for the classification and regression tasks, respectively.
@@
We apply post-fix masking to generate a binary mask ${\bm{m}}\in\{0,1\}^{V_{S}\cdot T^{\prime}}$ for the forecasting task. The encoder $f(\cdot)$ is used to compute the visible token features ${\bm{H}}\in\mathbb{R}^{N_{1}\times D}$. We then concatenate the sequence with learnable mask tokens to form ${\bm{H}}^{\prime}\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$, which is passed through the decoder $g(\cdot)$ to produce the final output. We optimise the MSE loss together with the NCC loss term over all reconstructed input tokens.
@@ 4.1 Model Variants and Implementation Details
We introduce OTiS in three different configurations, Base, Large, and Huge, with their specific architectures described in Appendix C.1, to explore scaling laws with respect to the model size. We set the patch size and stride to $P=24$, respectively, to split the time series into $T^{\prime}=\frac{\overline{T}}{P}$ non-overlapping patches along the time dimension. For pre-training, the context length specified in time points is set to $\overline{T}=1008$, resulting in $T^{\prime}=42$ sinusoidal temporal embeddings. If longer context lengths are required during fine-tuning, these embeddings are linearly interpolated (i.e. $T^{\prime}_{\text{ft}}\geq 42$) to offer greater flexibility for downstream applications. We tune the hyperparameters for pre-training and fine-tuning as described in Appendix C. An overview of the computational costs is provided in Appendix D.
@@ 4.2 Large and Diverse Pre-Training Corpus
We aim to develop a general time series model that fully handles the heterogeneity in real-world data. Specifically, our model is designed to handle time series with different variate counts $V_{S}$, inter-variate relationships, temporal dynamics, and sampling frequency, ensuring flexibility for downstream tasks. To this end, we pre-train our model on a large and diverse corpus of publicly available data spanning $8$ domains, with a total of $640,187$ samples and $11$ billion time points, as summarised in Table 1. A detailed description of the datasets included in our pre-training corpus can be found in Appendix A. The time series corpus is split into $612,394$ training and $27,793$ validation samples for pre-training.
Table 1: Overview of our large and diverse pre-training corpus. The corpus is built with unlabelled data from eight domains, encompassing medicine, engineering, natural sciences, and finance.
Domain Name Samples Variates Time points Frequency Disk size
$S$ $V_{S}$
ECG MIMIC-IV-ECG 2023 $400,000$ $12$ $5,000$ 500500500\,500Hz 909090\,90GB
Temperature DWD 2024 $203,340$ $1$ $720$ (hourly) 278278278\,278mHz 614614614\,614MB
Audio (stereo) AudioSet-$20$K 2017 $16,123$ $2$ $441,000$ 44.144.144.1\,44.1kHz 535353\,53GB
Audio (mono) AudioSet-$20$K 2017 $3,491$ $1$ $441,000$ 44.144.144.1\,44.1kHz 666\,6GB
Electromechanics FD-A 2016 $13,640$ $1$ $5,120$ 646464\,64kHz 161161161\,161MB
EEG TDBrain 2022 $2,692$ $19$ $60,000$ 500500500\,500Hz 121212\,12GB
EEG SEED 2015 $675$ $19$ $37,000$ 200200200\,200Hz 222\,2GB
Banking NN5555 2012 $111$ $1$ $971$ (daily) $12\,\mu$Hz 370370370\,370KB
Economics FRED-MD 2016 $107$ $1$ $728$ (monthly) 386386386\,386nHz 330330330\,330KB
Economics Exchange 2018 $8$ $1$ $7,588$ (daily) $12\,\mu$Hz 240240240\,240KB
640,187 11,052,756,981 164 GB
@@ 4.3 Benchmarking Across Domains and Tasks
To assess the utility of OTiS in real-world settings, we conduct experiments on three key use cases in time series analysis: classification, regression, and forecasting. For classification, we perform binary epilepsy detection using EEG (Epilepsy 2001), multi-class fault detection in rolling bearings from vibration signals (FD-B 2016), multi-class hand-gesture classification with accelerometer signals (Gesture 2009), and multi-class muscular disease classification using electromyographie (EMG 2000). In regression, we predict five imaging-derived cardiac phenotypes from $12$-lead ECG (LVEDV, LVESV, LVSV, LVEF, LVM 2020). For forecasting, we predict electricity transformer temperature (ETT 2021), weather (Weather 2024), and electricity consumption (Electricity 2024). Further dataset details are provided in Appendix B. We adhere to the established data splitting and evaluation procedures for the classification (Zhang et al., 2022), regression (Turgut et al., 2023), and forecasting (Zhou et al., 2021) tasks. All experiments are reported as mean and standard deviation across five seeds set during fine-tuning.
Table 2: Classification and regression performance on a total of $9$ benchmark tasks. OTiS is competitive with specialised baselines, setting new state-of-the-art on $6$ tasks and even outperforming the multimodal CM-AE and MMCL. This demonstrates the capability of OTiS to extract high-level semantics. Best score in bold, second best underlined. indicates tasks in previously unseen domains.
@@
* Models that incorporate paired imaging data during pre-training.
OTiS demonstrates strong classification capabilities, setting two new benchmarks, as shown in Table 2(a). Additionally, OTiS excels in predicting imaging-derived cardiac phenotypes, even surpassing multimodal baselines that incorporate imaging data during pre-training, as summarised in Table 2(b). Our model outperforms the baselines in $4$ out of $5$ regression tasks. As shown in Table 3, OTiS also exhibits strong forecasting capabilities, outperforming current models in $4$ out of $6$ benchmarks. Notably, all of the forecasting tasks are performed in previously unseen domains. A visualisation of the forecast predictions can be found in Appendix F. Overall, OTiS outperforms both specialised and general state-of-the-art baselines in $10$ out of $15$ diverse applications across $8$ domains, demonstrating its strong utility and generalisability across (unseen) domains and tasks.
Table 3: Forecasting performance on $6$ benchmark tasks. OTiS is competitive with specialised and general baselines, setting new state-of-the-art on $4$ tasks and showcasing its ability to capture local time series features. A forecasting horizon of $96$ time points is predicted from the past $336$ (*$512$, ${}^{+}904$) time points. Mean squared error (MSE $\downarrow$) is reported. Best score in bold, second best underlined. indicates tasks in previously unseen domains.

### v2

@@ 3.2 Pre-Training on Multi-Domain Time Series
To learn generalisable time series features, we adapt masked data modelling (MDM) (He et al., 2022) and pre-train OTiS on multi-domain time series, as visualised in Figure 2. We mask a subset of the input tokens and only encode the visible tokens using an encoder $f(\cdot)$. We then complement the encoded tokens with learnable mask tokens and pass them through a decoder $g(\cdot)$ to reconstruct the input tokens.
More precisely, we draw a binary mask $m\in\{0,1\}^{V_{S}\cdot T^{\prime}}$, following a dual masking strategy introduced in Section 3.4, and apply it to the input sequence $\bm{E}\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$. Thus, we obtain a visible view $\bm{E}[m]\in\mathbb{R}^{N_{1}\times D}$, where $N_{1}=\sum_{v=1}^{V_{S}}\sum_{t=1}^{T^{\prime}}m_{v,t}$ and $N_{0}=(V_{S}\cdot T^{\prime})-N_{1}$ denote the number of visible and masked tokens, respectively. The visible tokens $E[m]$ are then fed to the encoder $f(\cdot)$ to compute the time series features $\bm{H}\in\mathbb{R}^{N_{1}\times D}$:
$\bm{H}=f(\bm{E}[m]).$ (1)
To reconstruct the original input tokens, these features are passed through the decoder $g(\cdot)$ along with a learnable mask embedding $e^{\mathcal{M}}\in\mathbb{R}^{1\times D}$, which is inserted at each masked position where $m_{v,t}=0$:
$h^{\prime}_{v,t}=\begin{cases}h_{v,t}&\text{if }m_{v,t}=1\\ e^{\mathcal{M}}+e^{\mathcal{T}}_{t}+e^{\mathcal{V}}_{S,v}&\text{if }m_{v,t}=0\end{cases},$ (2)
such that $\bm{H}^{\prime}\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$. The decoder $g(\cdot)$ then predicts the reconstructed input $\bm{\widehat{X}}\in\mathbb{R}^{V_{S}\times(T^{\prime}\cdot P)}$:
$\bm{\widehat{X}}=g(\bm{H}^{\prime}),$ (3)
@@
$\mathcal{L}_{\text{MSE}}=\frac{1}{V_{S}\cdot T^{\prime}}\sum_{v=1}^{V_{S}}\sum_{t=1}^{T^{\prime}}\|x_{v,t}-\widehat{x}_{v,t}\|_{2}^{2}.$ (4)
@@ 3.3 Normalised Cross-Correlation Loss
MDM focuses on reconstructing masked parts of the data, emphasising local patterns through the MSE loss (4). However, time series often exhibit long-range dependencies, where past values influence future outcomes over extended periods. To accurately capture these global patterns, we introduce normalised cross-correlation (NCC) as a loss term
$\mathcal{L}_{\text{NCC}}=\frac{1}{V_{S}\cdot\overline{T}}\sum^{V_{S}}_{v=1}\sum^{\overline{T}}_{t=1}\hskip 2.84526pt\frac{1}{\sigma_{x_{v}}\sigma_{\widehat{x}_{v}}}(x_{v,t}-\mu_{x_{v}})(\widehat{x}_{v,t}-\mu_{\widehat{x}_{v}})$ (5)
for MDM in time series, where $\mu$ and $\sigma$ denote the mean and standard deviation, respectively, and $\mathcal{L}_{\text{NCC}}\in\left[-1,1\right]$. Hence, to capture both local and global temporal patterns, the total loss used to optimise OTiS is defined as
@@
where $\lambda\in\mathbb{R}$ is empirically set to $0.1$ during pre-training.
@@ 3.4 Dual Masking Strategy
We design the masking strategy to enhance foundational model capabilites in time series analysis. Specifically, we randomly select between two masking schemes during pre-training, namely random masking and post-fix masking. In 757575\,75of cases, we apply random masking, where each $m_{v,t}$ is independently sampled from a Bernoulli distribution with probability $p=1-\rho$, with $\rho\in[0,1]$ denoting the masking ratio (i.e. $m_{v,t}\sim\text{Bernoulli}(1-\rho)$). This encourages the model to learn complex inter-variate relationships across the entire time series. In the remaining $25\,\%$ of cases, we employ post-fix masking, which masks the second half of the temporal dimension, leaving only the first half visible (i.e. $m_{v,t}=\mathbbm{1}_{[t\leq T^{\prime}/2]}$). The prediction of future values solely based on past observations simulates real-world forecasting conditions, helping the model to capture temporal causality. Note that the optimal balance between the two masking schemes is empirically determined (see Appendix G.1). Overall, this dual masking strategy enables OTiS to learn both bidirectional relationships and temporal causality, which prove essential for general time series analysis.
3.5 Fine-Tuning & Inference
Adaptation to Unseen Domains.
For a previously unseen domain $S$, we randomly initialise variate embeddings $\bm{E}^{\mathcal{V}}_{S}\in\mathbb{R}^{V_{S}\times D}$. These embeddings are then fine-tuned along with the pre-trained encoder $f(\cdot)$ and, if required, the decoder $g(\cdot)$, for the specific task in $S$, as detailed in the following.
@@ 3.5 Fine-Tuning & Inference
We use the pre-trained encoder $f(\cdot)$ and the unmasked input sequence $\bm{E}$ to extract the time series features $\bm{H}=f(\bm{E})\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$. We average-pool these features into a global representation $h^{*}\in\mathbb{R}^{1\times D}$, which is passed through a linear layer to obtain the final model prediction. We optimise a cross-entropy and MSE loss for the classification and regression tasks, respectively.
@@
We use the pre-trained encoder $f(\cdot)$ and the unmasked input sequence $\bm{E}$ to extract the time series features $\bm{H}=f(\bm{E})\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$. We then concatenate the sequence with learnable mask embeddings $e^{\mathcal{M}}$ to form $\bm{H}^{\prime}\in\mathbb{R}^{(V_{S}\cdot(T^{\prime}+T^{\prime}_{H}))\times D}$, where $T^{\prime}_{H}\cdot P$ denotes the forecasting horizon specified in time points. This input sequence is then passed through the pre-trained decoder $g(\cdot)$ to obtain the final model prediction. We optimise both the MSE and NCC loss terms over all reconstructed input tokens.
@@ 4.1 Model Variants and Implementation Details
We introduce OTiS in three different configurations - Base, Large, and Huge (see Appendix C.1) - to explore scaling laws with respect to the model size. We set the patch size and stride to $P=24$, respectively, to split the time series into $T^{\prime}=\frac{\overline{T}}{P}$ non-overlapping patches along the temporal dimension. For pre-training, the context length is set to $\overline{T}=1008$ time points, resulting in $T^{\prime}=42$ sinusoidal temporal embeddings. If longer context lengths are required during fine-tuning, these embeddings are linearly interpolated (i.e. $T^{\prime}_{\text{ft}}\geq 42$) to offer greater flexibility for downstream applications. We tune the hyperparameters for pre-training and fine-tuning as described in Appendix C. An overview of the computational costs is provided in Appendix D.
Table 1: Overview of our pre-training corpus that spans eight domains from medicine, engineering, natural sciences, and finance.
Domain Name Samples Variates Time points Frequency
ECG MIMIC-IV-ECG $400,000$ $12$ $5,000$ 500500500\,500Hz
Temperature DWD $203,340$ $1$ $720$ $278\,\mu$Hz
Audio (stereo) AudioSet-$20$K $16,123$ $2$ $441,000$ 44.144.144.1\,44.1kHz
Audio (mono) AudioSet-$20$K $3,491$ $1$ $441,000$ 44.144.144.1\,44.1kHz
Electromechanics FD-A $13,640$ $1$ $5,120$ 646464\,64kHz
EEG TDBrain $2,692$ $19$ $60,000$ 500500500\,500Hz
EEG SEED $675$ $19$ $37,000$ 200200200\,200Hz
Banking NN5555 $111$ $1$ $971$ $12\,\mu$Hz
Economics FRED-MD $107$ $1$ $728$ 386386386\,386nHz
Economics Exchange $8$ $1$ $7,588$ $12\,\mu$Hz
640,187 11,052,756,981
@@ 4.2 Large and Diverse Pre-Training Corpus
We aim to develop a general time series model that fully handles the heterogeneity in real-world data. Specifically, our model is designed to handle time series with different variate counts $V_{S}$, inter-variate relationships, temporal patterns, and sampling frequency, ensuring flexibility for downstream tasks. To this end, we pre-train our model on a large and diverse corpus of publicly available data spanning $8$ domains, with a total of $640,187$ samples and $11$ billion time points, as summarised in Table 1. A detailed description of the datasets included in our pre-training corpus can be found in Appendix A. The time series corpus is split into $612,394$ training and $27,793$ validation samples for pre-training.
@@ 4.3 Benchmarking Across Domains and Tasks
We evaluate OTiS across general use cases in time series analysis: classification, regression, and forecasting. We use $12$ datasets from $8$ domains to compare our model against $27$ baselines, as outlined in Appendix B. The baselines comprise $15$ specialised, i.e. task-specific, models (either fully supervised or pre-trained and fine-tuned exclusively on target data) and $12$ general models (pre-trained on external data and fine-tuned on target data), including $6$ foundation models. We follow established data splitting and evaluation procedures for classification (Zhang et al., 2022), regression (Turgut et al., 2025), and forecasting (Zhou et al., 2021), with results reported across five seeds set during fine-tuning.
Table 2: Classification and regression results on a total of $9$ benchmark tasks, with indicating tasks in previously unseen domains. Best score in bold, second best underlined. OTiS successfully captures high-level semantics for classification and localised time series features for regression. Zero-shot and linear probing predictions indicate the generalisability of OTiS’ time series features.
@@
OTiS0 95.18 61.32 51.67 95.12
Zero-shot predictions of OTiS-Base.
@@
[…]
* Models incorporate paired imaging data during pre-training.
Linear probing of OTiS-Base.
Table 3: Forecasting results on $6$ benchmark tasks, with indicating tasks in previously unseen domains. A forecasting horizon of $96$ time points is predicted from the past $336$ (*$512$, ${}^{+}904$) time points. Mean squared error (MSE $\downarrow$) is reported. Best score in bold, second best underlined. OTiS successfully captures temporal patterns and trends essential for forecasting. Predictions involving minimal training () indicate the generalisability of OTiS’ extracted time series features.
@@
Model incorporates paired text data during pre-training and fine-tuning.
Predictions of OTiS-Base with only the domain-specific variate embeddings (VE) being fine-tuned.
The experiments reveal that OTiS extracts powerful time series features, achieving state-of-the-art performance on $10$ out of $15$ diverse benchmarks. The classification results in Table 2(a) highlight its particular strength in processing long time series, yielding a huge performance boost on FD-B ($\overline{T}_{\text{FD-B}}=5112$). The regression results in Table 2(b) emphasise the effectiveness of pre-training across domains, with OTiS outperforming specialised models that are unimodally pre-trained (e.g. MAE) or even multimodally pre-trained (e.g. MMCL). This is further supported by an additional ablation study on different training strategies presented in Appendix G.2. Moreover, the forecasting results in Table 3 and Appendix H reveal that OTiS effectively captures temporal patterns and trends, even in previously unseen domains. Finally, zero-shot experiments detailed in Appendix F, along with experiments involving minimal training (e.g. linear probing or fine-tuning of domain-specific embeddings) in Table 2(b) and 3, highlight how the extracted time series features generalise across domains and tasks, e.g. evidenced by the Epilepsy, LVEDV, and ETTh2 benchmarks.
(a) Frequency.
(b) Amplitude.
(c) Offset.
(d) Phase.
(e) Composition.
Figure 3: First two principal components of zero-shot sinusoidal representations extracted by OTiS-Base. We freeze OTiS after pre-training and use randomly initialised variate embeddings. The output features of the encoder are averaged to obtain a global representation. OTiS has an intrinsic understanding of fundamental time series properties, such as (a) frequency, (b) amplitude, (c) offset, and (d) phase. It effectively disentangles (e) simultaneous variations of these properties, providing a strong foundation for general time series analysis.
4.4 General Time Series Understanding
To evaluate whether OTiS has an intrinsic understanding of fundamental time series properties, such as frequency, amplitude, offset, and phase, we analyse zero-shot representations of sine waves, as illustrated in Figure 3. Particularly, we evaluate OTiS after pre-training without fine-tuning, using randomly initialised variate embeddings. We average the encoder’s output features to obtain a global representation.
When examining time series properties in isolation, we observe that their continuous variations are correctly mirrored in OTiS’ latent space. For frequency (see Figure 3(a)), the analysis reveals that embeddings for higher frequencies are positioned closer together, indicating greater similarity, than those for lower frequencies. This observation aligns with intuition: a change in frequency from 111\,1Hz to 505050\,50Hz represents a factor of $50$, thus resulting in more distinct embeddings, whereas a change from 505050\,50Hz to 100100100\,100Hz reflects only a doubling of the frequency. Analogous patterns are observed for amplitude (see Figure 3(b)) and offset (see Figure 3(c)). In the case of phase (see Figure 3(d)), the latent space forms a closed circle, with embeddings for phases of 00 and $2\pi$ being identical and positioned opposite to the embedding for a phase of $\pi$. These findings are consistent with the mathematical definition of a unit circle in the complex plane:
$e^{i\phi}=\text{cos}(\phi)+i\hskip 1.0pt\text{sin}(\phi),$ (7)
where $\phi\in[0,2\pi]$ denotes the phase. A phase of $\phi=2\pi$ corresponds to the identity (i.e. $e^{i2\pi}=e^{i0}$), while $\phi=\pi$ corresponds to a multiplication by $-1$ (i.e. $e^{i\pi}=-e^{i0}$).
Moreover, when faced with simultaneous variations in multiple time series properties (see Figure 3(e)), we observe that our model successfully disentangles them in its latent space. Overall, these findings demonstrate an intrinsic understanding of fundamental time series properties, providing a strong and essential foundation for general time series analysis.